# Florence, Italy - Pisa, Italy

### v1

@@ Abstract
Vision-Language Models (VLMs) demonstrate remarkable zero-shot generalization to unseen tasks, but fall short of the performance of supervised methods in generalizing to downstream tasks with limited data. Prompt learning is emerging as a parameter-efficient method for adapting VLMs, but state-of-the-art approaches require annotated samples. In this paper we propose a novel approach to prompt learning based on unsupervised knowledge distillation from more powerful models. Our approach, which we call Knowledge Distillation Prompt Learning (KDPL), can be integrated into existing prompt learning techniques and eliminates the need for labeled examples during adaptation. Our experiments on more than ten standard benchmark datasets demonstrate that KDPL is very effective at improving generalization of learned prompts for zero-shot domain generalization, zero-shot cross-dataset generalization, and zero-shot base-to-novel class generalization problems. KDPL requires no ground-truth labels for adaptation, and moreover we show that even in the absence of any knowledge of training class names it can be used to effectively transfer knowledge. The code is publicly available at https://github.com/miccunifi/KDPL. .
@@ 1 Introduction
Vision-Language Models (VLMs) like Contrastive Language-Image Pre-training (CLIP) are remarkably effective at zero-shot generalization to downstream tasks [42, 63, 2, 40]. These models leverage a dual encoder architecture and are trained to align image and text features in a shared embedding space.
CLIP and similar models are able to perform zero-shot classification by predicting the output class based on the similarity between the test image embedding and text embeddings of words from a fixed vocabulary.
@@
An intuitive way to achieve better performance is to use larger and more powerful VLMs [9] with better zero-shot generalization. However, though performance improves with increasing size, their practical applicability in real-world scenarios decreases.
For example, lightweight CLIP models based on ResNet-50 and ViT-B/32 require about 18 and 14 GFLOPs for inference, respectively. In contrast, more powerful models based on ViT-H/14 [17, 9] require nearly 30×30$\times$30 × the computation at 381 GFLOPs [17] for inference.
To address these challenges, parameter-efficient prompt learning is emerging as a promising and non-destructive approach to adapting VLMs to downstream tasks [62, 33, 34]. Zhou et al. [65] proposed a text-only prompt learning approach called CoOp to overcome the limitations of manually-crafted prompts. Subsequently, Jia et al. [28] extended prompt learning to the visual domain with Visual Prompt Tuning (VPT), and Khattaki et al. [29] proposed a multi-modal prompt learning approach called MaPLe. Although significantly improving over carefully tuned, hand-crafted prompts, these state-of-the-art techniques all require annotated samples from the dataset and have problems generalizing to other datasets or classes unseen during training [64, 29].
To eliminate the need for labeled training examples and improve the generalization of learned prompts, we propose a novel approach to prompt learning which we call Knowledge Distillation Prompt Learning (KDPL). KDPL adapts lightweight VLMs and improves performance on downstream tasks by distilling knowledge from a more powerful VLM without the need for annotated examples. KDPL is a flexible method that can be seamlessly integrated into existing prompt learning framework such as CoOp [65], CoCoOp [64], MaPLe [29], and VPT [28]. Figure 1 summarizes the motivations and illustrates the workflow of the proposed approach. We validate KDPL on two increasingly challenging scenarios: in the first, we assume knowledge of training class names, while in the second we assume no knowledge of training classes. Importantly, in both scenarios no annotated training samples are used.
Our contributions are:
@@ 2 Related Work
Prompt Learning. The performance of VLMs is highly dependent on textual prompt used to characterize downstream tasks. Given the significant impact of small changes in wording on performance, the manual generation of optimal textual prompts is a challenging task. Inspired by developments in NLP [62, 33, 34], CoOp [65] proposes to learn continuous vectors in the word embedding space instead of using tailored hand-crafted prompts.
The promising results of CoOp have attracted considerable interest in prompt learning for VLMs [35, 29, 55, 1, 6, 7, 30, 47]. CoCoOp, in particular, highlights the poor performance of CoOp on unseen classes and uses image-conditional dynamic prompts to improve generalization [64]. PLOT [7] instead is based on learning multiple prompts by exploiting local visual features. KgCoOp [55] adds a regularization loss to minimize the discrepancy between learned and hand-crafted prompts, while VPT [28] adapts prompt learning to the visual domain by learning continuous vectors in the input space of a ViT [15]. Finally, MaPLe [29] introduces a multimodal prompt learning strategy for VLMs. By tightly coupling vision-language prompts, MaPLe promotes synergistic learning while discouraging the collapse to independent unimodal solutions.
@@
3 Knowledge Distillation Prompt Learning
@@ 3.1 Preliminaries
Our approach is based on knowledge distillation applied to prompt learning. Here we discuss the base teacher and student models (CLIPs) and four state-of-the-art prompt learning approaches into which we will incorporate KDPL.
@@
To perform zero-shot classification using CLIP, we start with an image $I$ and build a set of textual prompts $\{Y_{i}\}_{i=1}^{C}$, where $C$ denotes the number of classes. Each handcrafted text prompt $Y_{i}$ takes the format “a photo of a [$\text{CLASS}_{i}$]”, where $\text{CLASS}_{i}$ represents a specific class name, such as airplane, bird, etc. Then the feature representations $\psi_{I}=f_{\theta}(I)$ and $\psi_{T}^{i}=g_{\phi}(E_{L}(Y_{i}))$ are extracted using the CLIP encoders. The predicted probability for each class is:
@@
where $\cos(\cdot,\cdot)$ denotes the cosine similarity and $\tau$ is a temperature hyperparameter.
@@
  • CoCoOp [64] extends CoOp by incorporating a lightweight network $h_{\theta}$. Each context token is obtained as $v_{i}(I)=v_{i}+\pi$, where $\pi=h_{\theta}(I)$. This method ensures that the textual prompts are conditioned on the input image as well.
@@
Note that methods involving visual prompt learning like VPT [28] and MaPLe [29] can only be applied to VLMs equipped with a ViT-based image encoder.
@@ 3.2 Label Agnostic Prompt Learning
The methods described above all rely on ground-truth labels during adaptation. Here we show how unsupervised knowledge distillation can be used instead.
@@
Prompt Learning via Unsupervised Knowledge Distillation. Given a lightweight CLIP model (the student) and a larger, more powerful CLIP model (the teacher), we aim to improve the downstream performance of the student model by distilling knowledge from teacher to student. For an image $I$ and a set of classes $\mathcal{C}=\{\text{CLASS}_{i}\}_{i=1}^{C}$, we start by performing zero-shot classification using the frozen teacher model. Specifically, we use the teacher image encoder $f_{\theta}^{B}$ and text encoder $g_{\phi}^{B}$ to compute the teacher image features $\psi_{I,B}=f_{\theta}^{B}(I)$ and text features $\psi_{T,B}^{i}=g_{\phi}^{S}(E_{L}([\text{CLASS}_{i}]))$. For the teacher model we use the fixed text prompt “a photo of [CLASS]”. We then apply Eq. 1 to produce the probabilities $p_{T}(I,\mathcal{C})$ predicted by the teacher on image $I$ for classes $\mathcal{C}$.
The teacher model does not rely on a learnable prompt and its predictions remain fixed during training. Our aim is to learn text and image prompts for the student model that enhance its generalization to downstream tasks. We denote with $f_{\theta}^{S}$ and $g_{\phi}^{S}$ the student image and text encoders, respectively, and with $\gamma$ the parameters associated with the learnable student prompts (see Fig. 2). Given the same image $I$ processed by the teacher and the same set of classes $\mathcal{C}$, the student extracts image features $\psi_{I,S}=f_{\theta,\gamma}^{S}(I)$ and text features $\psi_{T,S}^{i}=g_{\phi,\gamma}^{S}(E_{L}([\text{CLASS}_{i}]))$. Note that the text and image encoders can both depend on the prompt parameters $\gamma$. According to the prompt learning technique used, the $\gamma$ parameters can be used only by the text encoder (CoOp, CoCoOp), the visual encoder (VPT), or both (MaPLe). Finally, using Eq. 1 we produce student class probabilities $p_{S}(I,\mathcal{C})$ predicted on image $I$ for classes $\mathcal{C}$. Note that all encoder parameters except for the learnable prompt $\gamma$ are frozen during training.
@@
Importantly, our distillation loss does not assume any knowledge of class labels for image $I$. Nor does it require that $\mathcal{C}$ be the classes of the downstream task – that is, KDPL can be used for Label Agnostic and Class Agnostic adaptation scenarios. We found in early experiments that the symmetric KL-divergence works slightly better than either asymmetric option (see Section 4.6 for an ablation study on this choice).
@@ 3.3 Class Agnostic Prompt Learning
To further evaluate the generalization capabilities KDPL, we introduce a scenario where not only do we not know the labels of training images (label agnostic, Section 3.2) but where we also do not even know the class names associated with the dataset (class agnostic). This scenario is considerably more challenging as we make no assumptions about the few-shot training data.
@@ 4 Experimental Results
In this section we report on experiments performed to validate our proposed approach.
@@ 4.1 Evaluated Scenarios and Implementation Details
  • Class Agnostic Adaptation. We introduce a novel evaluation setting in which the training class names are unknown. The prompt is trained on ImageNet and tested on all the benchmark datasets used in the domain generalization and cross-dataset evaluations.
Implementation Details. We use a CLIP model with a ViT-H-14 visual backbone as the teacher model. [^2] https://huggingface.co/apple/DFN5B-CLIP-ViT-H-14 For student models, we evaluate both on a CLIP model based on ResNet-50 and a model based on ViT-B/32. By experimenting with both ResNet-based and ViT-based CLIP models we show the architecture independence of our approach.
To assess how KDPL performs when integrated into different prompt learning techniques, we selected four distinct textual, visual, and multimodal approaches.
With the ResNet-50 student model, we experiment with CoOp [65] and CoCoOp [64]. For the ViT-B/32 student, since it supports both visual and textual prompting, we experiment with CoOp [65], VPT [28], and MaPLe [29]. We denote the integration of our unsupervised knowledge distillation strategy with a specific prompt learning technique by adding the suffix +KDPL to the name of the corresponding technique. For example, CoOp+KDPL refers to the application of KDPL to CoOp. For class agnostic settings we instead use the suffix +CA-KDPL to indicate that no training class names are used.
@@
For each prompt learning method we use the original implementation and hyperparameters reported in the respective papers. Since we use the compact ResNet-50 and ViT-B/32 backbones, we cannot directly compare to the originally published results. Thus, all numbers we report here were computed by us and are averages over three independent runs. See the Supplementary Material for additional implementation details.
@@ 4.2 Results on Domain Generalization
Table 1 outlines the performance in the domain generalization setting. We report the performance of each baseline method alongside their unsupervised KDPL variants. Additionally, the performance of the zero-shot CLIP student and teacher models using a handcrafted prompt is included for comparison. We observe that applying our unsupervised approach to each baseline does not result in a decrease in performance on the source dataset. Notably, when transferring the learned prompts to a different domain, incorporating KDPL in each baseline leads to a slight improvement. This suggests that our unsupervised teacher-student distillation learns prompts that generalize better than those trained using ground-truth labels. The only exception where we observe an average decrease compared to the baseline is when using CoCoOp. In this case, CoCoOp-KDPL achieves lower average accuracy due to a significant gap on the ImageNet-R dataset. Finally, note that all the proposed unsupervised KDPL variants significantly improve the performance over the zero-shot CLIP student model in both the source and the target datasets.
@@ 4.3 Results on Cross-dataset Transfer
In Table 2 we present the results of the cross-dataset evaluation setting in which the prompt is trained on ImageNet and tested on ten different target datasets. For all datasets our KDPL-based variants consistently outperform the corresponding baselines and demonstrate superior generalization performance. The greater generalization capabilities of KDPL are evident even for fine-grained datasets like EuroSAT, on which CoOp+KDPL achieves an 8% improvement over the baseline CoOp when using ResNet-50 as the backbone. Although adding KDPL yields only minor improvement to the VPT and MaPLe prompt learning techniques, we emphasize that our VPT+KDPL and MaPLe+KDPL do not have access to ground-truth labels.

### v2

@@ Abstract
Vision-Language Models (VLMs) demonstrate remarkable zero-shot generalization to unseen tasks, but fall short of the performance of supervised methods in generalizing to downstream tasks with limited data. Prompt learning is emerging as a parameter-efficient method for adapting VLMs, but state-of-the-art approaches require annotated samples. In this paper we propose a novel approach to prompt learning based on unsupervised knowledge distillation from more powerful models. Our approach, which we call Knowledge Distillation Prompt Learning (KDPL), can be integrated into existing prompt learning techniques and eliminates the need for labeled examples during adaptation. Our experiments on more than ten standard benchmark datasets demonstrate that KDPL is very effective at improving generalization of learned prompts for zero-shot domain generalization, zero-shot cross-dataset generalization, and zero-shot base-to-novel class generalization problems. KDPL requires no ground-truth labels for adaptation, and moreover we show that even in the absence of any knowledge of training class names (CA-KDPL) can be used to effectively transfer knowledge. The code is publicly available at https://github.com/miccunifi/KDPL. .
@@ 1 Introduction
Vision-Language Models (VLMs) like Contrastive Language-Image Pre-training (CLIP) are remarkably effective at zero-shot generalization to downstream tasks [44, 66, 3, 42]. These models leverage a dual encoder architecture and are trained to align image and text features in a shared embedding space. CLIP and similar models are able to perform zero-shot classification by predicting the output class based on the similarity between the test image embedding and text embeddings of words from a fixed vocabulary.
@@
An intuitive way to achieve better performance is to use larger and more powerful VLMs [10] with better zero-shot generalization. However, though performance improves with increasing size, their practical applicability in real-world scenarios decreases. Lightweight CLIP models based on ResNet-50 and ViT-B/32 require about 18 and 14 GFLOPs for inference, respectively. In contrast, more powerful models based on ViT-H/14 [18, 10] require nearly 30×30$\times$30 × the computation at 381 GFLOPs [18] for inference.
To address these challenges, parameter-efficient prompt learning is emerging as a promising and non-destructive approach to adapting VLMs to downstream tasks [65, 34, 35]. Zhou et al. [68] proposed a text-only prompt learning approach called CoOp to overcome the limitations of manually crafted prompts. Subsequently, Jia et al. [29] extended prompt learning to the visual domain with Visual Prompt Tuning (VPT), and Khattak et al. [30, 31] proposed a multi-modal prompt learning approach called MaPLe [30] and a self-regulated approach called PromptSRC [31]. Although significantly improving over carefully tuned, hand-crafted prompts, these state-of-the-art techniques all require annotated samples from the dataset and have problems generalizing to other datasets or classes unseen during training [67, 30].
To eliminate the need for labeled training examples and improve the generalization of learned prompts, we propose a novel approach to prompt learning which we call Knowledge Distillation Prompt Learning (KDPL). KDPL adapts lightweight VLMs and improves performance on downstream tasks by distilling knowledge from a more powerful VLM without the need for annotated examples. KDPL is a flexible method that can be seamlessly integrated into existing prompt learning framework such as CoOp [68], CoCoOp [67], VPT [29], MaPLe [30], and PromptSRC [31]. Figure 1 summarizes the motivations and illustrates the workflow of the proposed approach. We validate KDPL on two increasingly challenging scenarios: in the first, we assume knowledge of training class names, while in the second we assume no knowledge of training classes. Importantly, in both scenarios no annotated training samples are used.
The contributions of this work are:
@@ 2 Related Work
Prompt Learning. The performance of VLMs is highly dependent on textual prompts used to characterize downstream tasks. Given the significant impact of small changes in wording on performance, the manual generation of optimal textual prompts is a challenging task. Inspired by developments in NLP [65, 34, 35], CoOp [68] proposes to learn continuous vectors in the word embedding space instead of using tailored hand-crafted prompts.
The promising results of CoOp have attracted considerable interest in prompt learning for VLMs [37, 30, 58, 2, 7, 8, 31, 49, 69, 50, 1]. CoCoOp, in particular, highlights the poor performance of CoOp on unseen classes and uses image-conditional dynamic prompts to improve generalization [67]. PLOT [8] instead is based on learning multiple prompts by exploiting local visual features. KgCoOp [58] adds a regularization loss to minimize the discrepancy between learned and hand-crafted prompts, while VPT [29] adapts prompt learning to the visual domain by learning continuous vectors in the input space of a ViT [16]. Finally, MaPLe [30] introduces a multimodal prompt learning strategy for VLMs. PromptSRC [31] introduces a self-regularization framework for prompting, addressing overfitting through agreement with a frozen model, a self-ensemble of prompts, and textual diversity.
@@
Techniques Using Teacher-Student Distillation. Recently Li et al. [36] introduced PromptKD which also uses teacher-student distillation in prompt learning. Although it too leverages a larger teacher to guide a smaller student, it differs from our contribution in several key aspects. PromptKD uses PromptSRC [31] to pre-train the teacher model with labeled examples, whereas KDPL does not pre-train the teacher and requires no labeled examples. PromptKD uses an MLP to project the student image encoder output, which adds extra parameters. KDPL is instead a general framework with no additional parameters, adaptable to any prompt-tuning technique to render them completely unsupervised. Moreover, PromptKD uses the entire training set, including both base and novel images, for distillation. In contrast, KDPL follows the standard 16-shot setting.
3 Knowledge Distillation Prompt Learning (KDPL)
@@ 3.1 Preliminaries
Our approach is based on knowledge distillation applied to prompt learning. Here we discuss the base teacher and student models (CLIPs) and five state-of-the-art prompt learning approaches into which we will incorporate KDPL.
@@
To perform zero-shot classification using CLIP, we start with an image $I$ and build a set of textual prompts $\{Y_{i}\}_{i=1}^{C}$, where $C$ denotes the number of classes. Each handcrafted text prompt $Y_{i}$ takes the format “a photo of a [$\text{CLASS}_{i}$]”, where $\text{CLASS}_{i}$ represents a specific class name, such as airplane, bird, etc.
Then the feature representations $\psi_{I}=f_{\theta}(I)$ and $\psi_{T}^{i}=g_{\phi}(E_{L}(Y_{i}))$ are extracted using the CLIP encoders. The predicted probability for each class is:
@@
where $\cos(\cdot,\cdot)$ is the cosine similarity and $\tau$ is a temperature hyperparameter.
@@
  • CoCoOp [67] extends CoOp by incorporating a lightweight network $h_{\theta}$. Each context token is obtained as $v_{i}(I)=v_{i}+\pi$, where $\pi=h_{\theta}(I)$. This method ensures that the textual prompts are conditioned on the input image.
@@
  • PromptSRC [31] is another multi-modal prompt learning technique that optimizes prompts through mutual agreement maximization between prompted and frozen model features using self-distillation, Gaussian-weighted prompt aggregation over the training session, and promoting textual diversity to address sample diversity imbalance.
Note that methods involving visual prompt learnable tokens like VPT [29], MaPLe [30], and PromptSRC [31] can only be applied to VLMs equipped with a ViT-based image encoder.
@@ 3.2 Label Agnostic Prompt Learning
The methods described above all rely on ground-truth labels during adaptation. Here we show how unsupervised knowledge distillation can be used instead to replace the need for annotated training examples.
@@
Prompt Learning via Unsupervised Knowledge Distillation (KDPL). Given a lightweight CLIP model (the student) and a larger, more powerful CLIP model (the teacher), we aim to improve the downstream performance of the student model by distilling knowledge from teacher to student. For an image $I$ and a set of classes $\mathcal{C}=\{\text{CLASS}_{i}\}_{i=1}^{C}$, we start by performing zero-shot classification using the frozen teacher model. Specifically, we use the teacher image encoder $f_{\theta}^{B}$ and text encoder $g_{\phi}^{B}$ to compute the teacher image features $\psi_{I,B}=f_{\theta}^{B}(I)$ and text features $\psi_{T,B}^{i}=g_{\phi}^{S}(E_{L}([\text{CLASS}_{i}]))$. For the teacher model we use the fixed text prompt “a photo of [CLASS]”. We then apply Eq. 1 to produce the probabilities $p_{T}(I,\mathcal{C})$ predicted by the teacher on image $I$ for classes $\mathcal{C}$.
The teacher model does not rely on a learnable prompt and its predictions remain fixed during training. Our aim is to learn text and image prompts for the student model that enhance its generalization to downstream tasks.
We denote with $f_{\theta}^{S}$ and $g_{\phi}^{S}$ the student image and text encoders, respectively, and with $\gamma$ the parameters associated with the learnable student prompts (see Fig. 2). Given the same image $I$ processed by the teacher and the same set of classes $\mathcal{C}$, the student extracts image features $\psi_{I,S}=f_{\theta,\gamma}^{S}(I)$ and text features $\psi_{T,S}^{i}=g_{\phi,\gamma}^{S}(E_{L}([\text{CLASS}_{i}]))$. Note that the text and image encoders can both depend on the prompt parameters $\gamma$. According to the prompt learning technique used, the $\gamma$ parameters can be used only by the text encoder (CoOp, CoCoOp), the visual encoder (VPT), or both (MaPLe, PromptSRC). Finally, using Eq. 1 we produce student class probabilities $p_{S}(I,\mathcal{C})$ predicted on image $I$ for classes $\mathcal{C}$. Note that all encoder parameters except for the learnable prompt $\gamma$ are frozen.
@@
Importantly, our distillation loss does not assume any knowledge of class labels for image $I$. Nor does it require that $\mathcal{C}$ be the classes of the downstream task – that is, KDPL can be used for Label Agnostic and Class Agnostic adaptation scenarios. We found in early experiments that the symmetric KL-divergence works slightly better than either asymmetric option (see Section B.3. Symmetric versus Asymmetric KL in the Supplementary Material for an ablation study on this choice).
@@ 3.3 Class Agnostic Prompt Learning
To further evaluate the generalization capabilities of KDPL, we introduce a scenario where not only do we not know the labels of training images (label agnostic, Section 3.2) but where we also do not even know the class names associated with the dataset (class agnostic). This scenario is considerably more challenging as we make no assumptions about the few-shot training data.
@@ 4 Experimental Results
In this section we report on experiments validating our proposed approach.
@@ 4.1 Evaluated Scenarios and Implementation Details
  • Class Agnostic Adaptation. In addition we introduce a novel evaluation setting in which the training class names are unknown. The prompt is trained on ImageNet and tested on all the benchmark datasets used in the domain generalization and cross-dataset evaluations.
Implementation Details. We use a CLIP model with a ViT-H-14 visual backbone as the teacher model. [^2] https://huggingface.co/apple/DFN5B-CLIP-ViT-H-14 For student models, we evaluate both on a CLIP model based on ResNet-50 and a model based on ViT-B/32. By experimenting with both ResNet-based and ViT-based CLIP models we show the architecture independence of our approach. To assess how KDPL performs when integrated into different prompt learning techniques, we selected five distinct textual, visual, and multimodal approaches. With the ResNet-50 student model, we experiment with CoOp [68] and CoCoOp [67]. For the ViT-B/32 student, since it supports both visual and textual prompting, we experiment with CoOp [68], VPT [29], MaPLe [30] and PromptSRC [31].
We denote the integration of our unsupervised knowledge distillation strategy with a specific prompt learning technique by adding the suffix +KDPL to the name of the corresponding technique. For example, CoOp+KDPL refers to the application of KDPL to CoOp. For class agnostic settings we instead use the suffix +CA-KDPL to indicate that no training class names are used.
@@
For each prompt learning method we use the original implementation and hyperparameters reported in the respective papers. Since we use the compact ResNet-50 and ViT-B/32 backbones, we cannot directly compare to the originally published results. Thus, all numbers we report here were computed by us and are averages over three independent runs. See Section A. Additional Implementation Details in the Supplementary Material for additional implementation details.
@@ 4.2 Results on Domain Generalization
Table 1 outlines the performance in the domain generalization setting. We report the performance of each baseline method alongside their unsupervised KDPL variants. Additionally, the performance of the zero-shot CLIP student and teacher models using a handcrafted prompt is included for comparison. We observe that applying our unsupervised approach to each baseline does not result in a significant decrease in performance on the source dataset. Notably, when transferring the learned prompts to a different domain, incorporating KDPL in each baseline can lead to a slight improvement. This suggests that our unsupervised teacher-student distillation learns prompts that generalize better than those trained using ground-truth labels. The only exceptions where we observe an average decrease compared to the baseline are when using CoCoOp or PromptSRC. CoCoOp-KDPL achieves an average accuracy significantly lower than CoCoOp only on the ImageNet-R dataset. Finally, note that all our unsupervised KDPL variants significantly improve the performance over the zero-shot CLIP student model in both the source and the target datasets.
@@ 4.3 Results on Cross-dataset Transfer
In Table 2 we present the results of the cross-dataset evaluation setting in which the prompt is trained on ImageNet and tested on ten different target datasets. For all datasets our KDPL-based variants consistently outperform the corresponding baselines and demonstrate superior generalization performance. The greater generalization capabilities of KDPL are evident even for fine-grained datasets like EuroSAT, on which the CoOp+KDPL achieves an 8% improvement over the baseline CoOp when using ResNet-50 as the backbone. Although adding KDPL yields only minor improvement to the VPT and MaPLe prompt learning techniques, we emphasize that our VPT+KDPL and MaPLe+KDPL do not have access to ground-truth labels. Notably, PromptSRC+KDPL achieves the highest average performance.